# A Smart Mnemonic Sounds like “Glue Tonic”: Mixing LLMs with Student Feedback to Make Mnemonic Learning Stick

### v1

@@ Abstract
Keyword mnemonics are memorable explanations that link new terms to simpler keywords. Prior works generate mnemonics for students, but they do not guide models toward mnemonics students prefer and aid learning. We build Smart, a mnemonic generator trained on feedback from real students learning new terms. To train Smart, we first fine-tune LLaMA-2 on a curated set of user-written mnemonics. We then use LLM alignment to enhance Smart: we deploy mnemonics generated by Smart in a flashcard app to find preferences on mnemonics students favor. We gather 2684 preferences from 45 students across two types: expressed (inferred from ratings) and observed (inferred from student learning), yielding three key findings. First, expressed and observed preferences disagree; what students think is helpful does not fully capture what is truly helpful. Second, Bayesian models can synthesize complementary data from multiple preference types into a single effectiveness signal. Smart is tuned via Direct Preference Optimization on this signal, which we show resolves ties and missing labels in the typical method of pairwise comparisons, augmenting data for LLM output quality gains. Third, mnemonic experts assess Smart as matching GPT-4, at much lower deployment costs, showing the utility of capturing diverse student feedback to align LLMs in education.111 https://github.com/nbalepur/Mnemonic
@@ 1 Mnemonics Aid Vocabulary Learning
Keyword mnemonics promote efficient and engaging vocabulary (vocab) learning Benge and Robbins (2009). These tools help students learn a new term’s meaning (e.g. Benevolent) by relating it to a simpler keyword (e.g. Benevolent sounds like benefit), and explaining how the keyword and term are linked (e.g. A boss giving employee benefits is kind, which is the meaning of benevolent) Pressley et al. (1982). Students use mnemonics to prepare for exams like the GRE Fairbanks (1977) which involve mastering hundreds of terms Al-Jarf (2016). Despite their utility, writing mnemonics is tedious, requiring vocabulary expertise and creativity to make memorable keyword links Siriganjanavong (2013).
@@
To ease these burdens, prior work has automatically generated keyword mnemonics Savva et al. (2014). However, most works test keyword extraction Anonthanasap et al. (2015), omitting the explanations linking keywords to terms that enable effective mnemonic use Raugh and Atkinson (1975). Large Language Models (LLMs) are apt for writing explanations, a difficult task that tests if LLMs can combine vocabulary Huang et al. (2022a), phonology Suvarna et al. (2024), commonsense Davis (2023), and creativity Tian et al. (2023) to help students learn (section 6.3). While promising, existing works only prompt LLMs Lee and Lan (2023) and lack training on student feedback to guide LLMs toward mnemonics students prefer and benefit learning.
@@
In pursuit of student-guided mnemonics, we propose Smart, which employs Student Mnemonic Alignment to generate keyword mnemonics that aid the Recall of vocab Terms. (Figure 1). To train Smart, we first collect data from MnemonicDictionary Memliapp (2007), a site where users submit mnemonics they find helpful. We curate a high-quality subset of this data and fine-tune LLaMA-2 70B (Figure 1, left) as our initial model (section 2). To enhance Smart, we draw from LLM alignment, which improves LLMs via tuning to preference labels that capture which of two LLM outputs users favor Casper et al. (2023). We gather preferences by sampling mnemonics from our initial model and deploying them to students in a flashcard app (section 3).
@@
There are many ways to collect preferences, and Bansal et al. (2024) show that pairwise rankings and Likert ratings yield conflicting labels on which LLM output is favored. To study preference agreement in education (Figure 1, mid), we also gather pairwise and Likert annotations, which we define as expressed preferences: those inferred from user ratings. Expressed preferences measure what users think are more helpful, but to see if this agrees with what truly helps users, we introduce observed preferences: those inferred from observable outcomes as users interact with outputs. We collect observed preferences via the mean time users need to learn a term while studying with its mnemonic, serving as a proxy for mnemonic short-term learning efficacy.
@@
Over three months, 45 students give us 2684 preferences on mnemonic pairs. To decide how to align Smart, we study the relation of preference types. Expressed and observed preferences disagree (section 4), so what students think helps them learn differs from what truly helps them learn. We intuit these preference types represent equally-valuable goals (section 5): an effective mnemonic should be non-harmful and liked by users (expressed), but also aid learning (observed). Thus, we design a Bayesian model Gelman and Hill (2006) that learns mnemonic effectiveness via feedback from all preference types (section 5.1). We compare mnemonics by effectiveness to elect a winning mnemonic in the pair (Figure 1, right) and tune Smart with Direct Preference Optimization (Rafailov et al., 2024, DPO) on this signal (section 5.2).
@@
We conduct several experiments to test Smart. Fine-tuning and DPO enhance Smart, so aligning LLMs to student preferences improves educational text (section 6.1). Further, combining all preferences via Bayesian modeling can resolve ties or missing labels in the typical method of pairwise comparisons, augmenting DPO data for mnemonic quality gains (section 6.2). Multiple preferences can be gathered in one app, sometimes with no extra annotations (section 3.2.2). So if resources allow, we advise collecting multiple preferences to study the relation of complementary alignment objectives, which can then be combined for data augmentation to improve LLM outputs.
@@
Lastly, two mnemonic experts assess mnemonics (section 6.3) from Smart, GPT-4, and an expert human writer, finding: 1) Smart matches the SOTA LLM GPT-4, showing the utility of student feedback; and 2) Our writer largely surpasses models at creating simple keywords and imageable explanations, motivating mnemonic generation as a difficult task and giving insights into feedback that can be used to guide LLMs in similar tasks. Our contributions are:
@@
1) We design Smart, an LLM mnemonic generator guided by feedback from real-world students.
2) We analyze expressed and observed preferences, finding that students cannot predict their learning.
3) We align Smart with multiple preferences via Bayesian modeling and DPO, resulting in keyword mnemonics that match the SOTA LLM GPT-4.
@@ 2.1 Data Collection
There does not exist a dataset of vocab terms and mnemonics, so we curate new datasets to facilitate mnemonic research. We use vocab words from the Graduate Records Examination (GRE), a graduate admissions exam that students prepare for by learning hundreds of vocab terms Nayak et al. (2017). Mnemonics have been used to help students learn GRE vocabulary Fairbanks (1977); Pi et al. (2021).
@@ 2.2 Identifying High-Quality Mnemonics
The user-submitted mnemonics collected from section 2.1 are noisy, but using a subset of high-quality data can better train Smart for mnemonic generation Xia et al. (2024). MnemonicDictionary users upvote or downvote mnemonics, so upvote ratio could find high-quality data, but this metric does not consider all upvotes given Powell et al. (2017). Thus, following Hoffart et al. (2019), we use a Bayesian model to learn the probability $q_{i}$ of mnemonic $m_{i}$ being high-quality, based on the upvote $\nu_{u,i}$ and downvote $\nu_{d,i}$ counts on $m_{i}$. We assume mnemonics with higher $q_{i}$ have more upvotes, so we model $\nu_{u,i}$ as a Binomial distribution with probability $q_{i}$:
$q_{i}$ $\sim\mathrm{Beta}(\alpha=2,\beta=8)$ (1)
$\nu_{u,i}$ $\sim\mathrm{Binomial}(\nu_{u,i}+\nu_{d,i},q_{i})$ (2)
$q_{i}$ has the prior $\alpha=2,\beta=8$, as our brief manual assessment found that ${\sim}20\%$ of the mnemonics are high-quality. We estimate $q_{i}$ via No U-Turn Sampling (Hoffman et al., 2014, NUTS). Pairs ($v_{i}$, $m_{i}$) with the 1000-highest $q_{i}$ values form the fine-tuning dataset $\mathcal{D}_{ft}$ for Smart (details in Appendix A).
@@ 2.3 Model Fine-Tuning
The dataset $\mathcal{D}_{ft}$ has term/mnemonic pairs $(v,m)$, so we can use $\mathcal{D}_{ft}$ to train an initial seq2seq Smart model $p_{0}(m|v)$ to generate $m$ from $v$. Upon inspection, we find some quality issues in the mnemonics, so we use GPT-4 to clean grammar errors in $m$ via a 0-shot prompt, and discard any $m$ with offensive or overly culturally-specific text (see Appendix A.1). We end up with 889 pairs for fine-tuning Smart.
[…]
$\mathcal{L}_{CE}=\sum_{j=1}^{|m|}\log p(m_{j}|m_{1},...,m_{j-1},\mathcal{P})$ (3)
@@ 3 Collecting Mnemonic Preferences
Only fine-tuning does not explicitly guide Smart toward mnemonics that users prefer and help them learn—our overall goal. Thus, we use alignment Ziegler et al. (2019), which tunes LLMs to preference labels capturing which outputs users favor. To align Smart, we need a preference dataset $\mathcal{D}_{pref}$ with entries of a term $v$, mnemonic pair $(m_{A},m_{B})$ for $v$ created by the initial model $p_{0}(m|v)$, and preference label $y\in\{\texttt{A},\texttt{B},\texttt{tie}\}$ noting the mnemonic in the pair users favor. To build $\mathcal{D}_{pref}$, we create mnemonic pairs (section 3.1), define our preference labels (section 3.2), and describe our user study details (section 3.3).
@@ 3.1 Generating Mnemonic Pairs
For $\mathcal{D}_{pref}$, we first need mnemonic pairs $(m_{A}$,$m_{B})$ created by model $p_{0}(m|v)$ for many terms $v$. For optimal results, we seek diverse mnemonics with high probability in $p_{0}(m|v)$. This ensures the pair reflects the distribution of $p_{0}(m|v)$ and gives the user distinct choices, yielding clearer preferences.
We do this via best-of-$n$ sampling Nakano et al. (2021), which samples $n$ LLM outputs and picks the one with the best score from a reward model. We define a reward $\pi_{pair}(m_{A},m_{B},v)$ that returns the sum of $m_{A}$ and $m_{B}$ sequence probabilities from $p_{0}(m|v)$, minus the ROUGE-1 Lin (2004) of $m_{A}$ and $m_{B}$. The reward favors mnemonics with high sequence probability and low word overlap. To create mnemonic pairs, we take 1000 terms $\mathcal{V}_{pref}\subset\mathcal{V}$ not used in $\mathcal{D}_{train}$. For each $v\in\mathcal{V}_{pref}$, we sample five mnemonics =$\mathcal{M}=\{m\}^{5}\sim p_{0}(m|v)$ with 0.3 temperature. We take $(m_{A},m_{B})\in\mathcal{M}\times\mathcal{M}$ with the best $\pi_{pair}(m_{A},m_{B},v)$ score as the pair for $v$.
@@ 3.2 Preference Label Collection
While preferences are often elicited through crowdworker sites, we decide to collect preferences from students who can directly benefit from mnemonics. Flashcard software can aid mnemonic use Tuite et al. (2012), so we host our user annotation schema within a flashcard app to gather student preferences $y\in\{\texttt{A},\texttt{B},\texttt{tie}\}$ for the mnemonic pairs in section 3.1.
In our app, users study flashcards $f_{v}$ with a term $v$ as the front of the card and type its definition (Figure 2). If a user answers $f_{v}$ incorrectly, they see a mnemonic from the pair (section 3.1) for $v$ to aid learning (Figure 3). Users keep studying with $f_{v}$ until they answer $f_{v}$ correctly (after $n$ turns). We use the $\text{KAR}^{\text{3}}\text{L}$ Shu et al. (2024) model to schedule cards.
@@ 3.2.1 Expressed Preferences
For Likert ratings, when a user sees a mnemonic $m_{X}$ after answering card $f_{v}$ incorrectly, they rate $m_{X}$ on a 5-Likert scale (Figure 3). We call this preference label $y_{rate}$ and set $y_{rate}=\texttt{A}$ if the average Likert rating of $m_{A}$ is higher than $m_{B}$ across users (same for B and tie). For each $f_{v}$, users see only one of $m_{A}$ or $m_{B}$, so their rating cannot be biased by having already seen the other in the pair.

### v2

@@ Abstract
Keyword mnemonics are memorable explanations that link new terms to simpler keywords. Prior work generates mnemonics for students, but they do not train models using mnemonics students prefer and aid learning. We build Smart, a mnemonic generator trained on feedback from real students learning new terms. To train Smart, we first fine-tune LLaMA-2 on a curated set of user-written mnemonics. We then use LLM alignment to enhance Smart: we deploy mnemonics generated by Smart in a flashcard app to find preferences on mnemonics students favor. We gather 2684 preferences from 45 students across two types: expressed (inferred from ratings) and observed (inferred from student learning), yielding three key findings. First, expressed and observed preferences disagree; what students think is helpful does not always capture what is truly helpful. Second, Bayesian models can synthesize complementary data from multiple preference types into a single effectiveness signal. Smart is tuned via Direct Preference Optimization on this signal, which resolves ties and missing labels in the typical method of pairwise comparisons, augmenting data for LLM output quality gains. Third, mnemonic experts assess Smart as matching GPT-4 at much lower deployment costs, showing the utility of capturing diverse student feedback to align LLMs in education.111 https://github.com/nbalepur/Mnemonic
@@ 1 Mnemonics Aid Vocabulary Learning
Keyword mnemonics promote efficient and engaging vocabulary (vocab) learning Benge and Robbins (2009). These tools help students learn a new term’s meaning (e.g. Benevolent) by relating it to a simpler keyword (e.g. Benevolent sounds like benefit), and explaining how the keyword and term are linked (e.g. A boss giving employee benefits is kind, which is the meaning of benevolent) Pressley et al. (1982). Students use mnemonics to prepare for exams like the GRE Fairbanks (1977) which involve mastering hundreds of terms Khan (2009). Despite their utility, writing mnemonics is tedious, requiring vocabulary expertise and creativity to make memorable keyword links Siriganjanavong (2013).
@@
To ease these burdens, prior work automatically generates keyword mnemonics Savva et al. (2014). However, most works design keyword extractors Anonthanasap et al. (2015), omitting the explanations linking keywords to terms that enable effective mnemonic use Raugh and Atkinson (1975). Large Language Models (LLMs) are apt for writing explanations, a difficult task that tests if LLMs can combine vocabulary Huang et al. (2022a), phonology Suvarna et al. (2024), commonsense Davis (2023), and creativity Tian et al. (2024) to help students learn (section 6.3). While promising, existing works only prompt LLMs Lee and Lan (2023) and lack training on student feedback to guide LLMs toward mnemonics students prefer and benefit learning.
@@
In pursuit of student-guided mnemonics, we propose Smart, which employs Student Mnemonic Alignment to generate keyword mnemonics that aid the Recall of Terms. (Figure 1). To train Smart, we first get data from MnemonicDictionary Memliapp (2007), a site where users submit mnemonics that they find helpful. We collect a high-quality subset of submitted mnemonics to fine-tune LLaMA-2 70B (Figure 1, left) as our initial model (section 2). To enhance Smart, we draw from LLM alignment, which improves LLMs via tuning to preference labels that capture which of two LLM outputs users favor Casper et al. (2023). We gather preferences by sampling mnemonics from our initial model and deploying them to students in a flashcard app (section 3).
@@
There are many ways to collect preferences, and Bansal et al. (2024) show that pairwise rankings and Likert ratings yield conflicting labels on which LLM output is favored. To study preference agreement in education (Figure 1, mid), we also gather pairwise and Likert annotations, which we define as expressed preferences: those inferred from user ratings. Expressed preferences measure what users think are more helpful, but to see if this agrees with what truly helps users, we introduce observed preferences: those inferred from observable goals (e.g. learning) as users interact with outputs. We collect observed preferences via the mean time users need to learn a term while studying with its mnemonic, a proxy for mnemonic short-term learning efficacy.
@@
Over three months, 45 students gave 2684 preferences on mnemonic pairs. To decide how to align Smart, we study the relation of preference types. Expressed and observed preferences disagree (section 4), so what students think helps them learn differs from what truly helps them learn. These preference types represent equally valuable goals (section 5): an effective and helpful mnemonic should be non-harmful and liked by users (expressed), but also aid learning (observed). Thus, we design a Bayesian model Gelman and Hill (2006) that learns mnemonic effectiveness via feedback from all preference types (section 5.1). We compare mnemonics by effectiveness to elect a winning mnemonic in the pair (Figure 1, right) and tune Smart with Direct Preference Optimization (Rafailov et al., 2024, DPO) on this signal (section 5.2).
@@
We assess Smart through several experiments. Fine-tuning and DPO enhance Smart, so aligning LLMs to student preferences improves educational text (section 6.1). Further, combining all preferences via Bayesian modeling can resolve ties or missing labels in the typical method of pairwise comparisons, augmenting DPO data for mnemonic quality gains (section 6.2). Multiple preferences can be gathered in one app, sometimes with no extra annotations (section 3.2.2). So if resources allow, we advise collecting multiple preferences to study the relation of complementary alignment objectives, which can then be combined for data augmentation to improve LLM outputs.
@@
Lastly, two mnemonic experts assess mnemonics (section 6.3) from Smart, GPT-4, and a freelance creative writer from Upwork, finding: 1) Smart matches the SOTA LLM GPT-4, showing the utility of student feedback; and 2) Our writer’s keywords are much simpler and their explanations are more imageable compared to GPT-4 and Smart, motivating mnemonic generation as a difficult task and giving insights into feedback types (simplicity, imageability) that can also be collected to better align LLMs for downstream tasks. Our contributions are:
@@
1) We design Smart, an LLM mnemonic generator aligned by feedback from real-world students.
2) We analyze expressed and observed preferences, finding that the LLM outputs students think help them learn is not what actually helps them learn.
3) We align Smart with multiple preferences via Bayesian modeling, which can break preference ties for DPO output quality gains and results in a more efficient keyword mnemonic generator that matches the state-of-the-art LLM GPT-4.
@@ 2.1 Data Collection
A dataset with vocab terms and mnemonics does not exist, so we curate new datasets to facilitate mnemonic research. We use vocab words from the Graduate Records Examination (GRE), a graduate admissions exam that students prepare for by learning hundreds of vocab terms Nayak et al. (2017). Mnemonics have been used to help students learn GRE vocabulary Fairbanks (1977); Pi et al. (2021).
@@ 2.2 Identifying High-Quality Mnemonics
The user-submitted mnemonics collected from section 2.1 are noisy, but a subset of high-quality data would better train Smart for mnemonic generation Xia et al. (2024). MnemonicDictionary users upvote or downvote mnemonics, so upvote ratio could find high-quality data, but this metric does not consider all upvotes given Powell et al. (2017). Thus, following Hoffart et al. (2019), we build a Bayesian model to learn the probability $q_{i}$ of mnemonic $m_{i}$ being high-quality, based on the upvote $\nu_{u,i}$ and downvote $\nu_{d,i}$ counts on $m_{i}$. We assume mnemonics with higher $q_{i}$ have more upvotes, so we model $\nu_{u,i}$ as a Binomial distribution with probability $q_{i}$:
$q_{i}$ $\sim\mathrm{Beta}(\alpha=2,\beta=8),$ (1)
$\nu_{u,i}$ $\sim\mathrm{Binomial}(\nu_{u,i}+\nu_{d,i},q_{i}),$ (2)
which has prior $\alpha=2,\beta=8$, as our brief manual assessment found that ${\sim}20\%$ of the mnemonics are high-quality. We estimate $q_{i}$ via No U-Turn Sampling (Hoffman et al., 2014, NUTS). Pairs ($v_{i}$, $m_{i}$) with the 1000-highest $q_{i}$ values form the fine-tuning dataset $\mathcal{D}_{ft}$ for Smart (details in Appendix A).
@@ 2.3 Model Fine-Tuning
The dataset $\mathcal{D}_{ft}$ has term/mnemonic pairs $(v,m)$, so we can use $\mathcal{D}_{ft}$ to train an initial seq2seq Smart model $p_{0}(m\,|\,v)$ to create $m$ from $v$. Upon inspection, we find some quality issues in the mnemonics, so we use GPT-4 to clean grammar errors in $m$ via a 0-shot prompt, and discard any $m$ with offensive or overly culturally-specific text (see Appendix A.1). We end up with 889 pairs for fine-tuning Smart.
Each $(v,m)\in\mathcal{D}_{ft}$ forms prompt $\mathcal{P}=\text{``}\texttt{Term:}$: $v\backslash\texttt{nMnemonic:}\text{''}$: ” and output text $m$. Our initial model $p_{0}(m\,|\,v)$ fine-tunes LLaMA-2 70B Touvron et al. (2023) to minimize the cross-entropy loss $\mathcal{L}_{CE}$ of predicting tokens $m_{j}\in m$ given $\mathcal{P}$:
$\mathcal{L}_{CE}=\sum_{j=1}^{|m|}\log p(m_{j}\,|\,m_{1},...,m_{j-1},\mathcal{P}).$ (3)
@@ 3 Collecting Mnemonic Preferences
Only fine-tuning does not explicitly guide Smart toward mnemonics that users prefer and help them learn—our overall goal. Thus, we use alignment Ziegler et al. (2019): tuning LLMs to preference labels capturing which outputs users favor. To align Smart, we need a preference dataset $\mathcal{D}_{pref}$ with entries of a term $v$, mnemonic pair $(m_{A},m_{B})$ for $v$ created by the initial model $p_{0}(m\,|\,v)$, and preference label $y\in\{\texttt{A},\texttt{B},\texttt{tie}\}$ noting the mnemonic in the pair users favor. To build $\mathcal{D}_{pref}$, we create mnemonic pairs (section 3.1), define our preference labels (section 3.2), and describe our user study details (section 3.3).
@@ 3.1 Generating Mnemonic Pairs
For $\mathcal{D}_{pref}$, we first need mnemonic pairs $(m_{A}$,$m_{B})$ created by model $p_{0}(m\,|\,v)$ for many terms $v$. In preference datasets, researchers sample candidate LLM responses to match the abilities of the target LLM Bai et al. (2022), so the mnemonics should have high probability in our model $p_{0}(m\,|\,v)$. Further, to ensure the mnemonics in the pair are not too similar, which would often result in preference ties, we seek a diverse pair of mnemonics; this gives the user distinct choices, yielding clearer preferences.
We combine these two objectives through best-of-$n$ sampling Nakano et al. (2021), which samples $n$ LLM outputs and picks the one with the best score from a reward model. We define a reward $\pi_{pair}(m_{A},m_{B},v)$ that returns the sum of $m_{A}$ and $m_{B}$ sequence probabilities from $p_{0}(m\,|\,v)$, minus the ROUGE-1 Lin (2004) of $m_{A}$ and $m_{B}$, assessing mnemonic diversity Shaib et al. (2024). The reward favors mnemonics with high sequence probability and low word overlap. To create mnemonic pairs, we sample 1000 terms $\mathcal{V}_{pref}\subset\mathcal{V}$ not used in $\mathcal{D}_{train}$. For each term $v\in\mathcal{V}_{pref}$, we sample five mnemonics =$\mathcal{M}=\{m\}^{5}\sim p_{0}(m\,|\,v)$ with 0.3 temperature. We take $(m_{A},m_{B})\in\mathcal{M}\times\mathcal{M}$ with the best $\pi_{pair}(m_{A},m_{B},v)$ score as the pair for $v$.
@@ 3.2 Preference Label Collection
While preferences are often elicited through crowdworker sites, we decide to collect preferences from students who can learn vocab using our mnemonics. Flashcard software can aid mnemonic use Tuite et al. (2012), so we host our user annotation schema within a flashcard app to gather student preferences $y\in\{\texttt{A},\texttt{B},\texttt{tie}\}$ for the mnemonic pairs in section 3.1.
Flashcards have two sides and while studying, users read the front of the card and answer what is on the back. In our app, users study flashcards $f_{v}$ with a term $v$ as the front of the card and type its definition on the back (Figure 2). In one session, users study a set of flashcards with 5 to 50 terms the user has not yet studied, and continue studying until they correctly type each term’s definition. We use TF-IDF Sparck Jones (1972) with a cutoff of 0.15 to check if the user’s typed definition matches the ground-truth definition, which the user can override if they disagree with the metric’s prediction. If a user answers $f_{v}$ correctly, it is removed from the pool of cards left to study. If answered incorrectly, they see a mnemonic from the pair (section 3.1) for $v$ to aid learning (Figure 3), and the card remains in the set of cards to study. Thus, for each card $f_{v}$, we can count how many turns $n\in\mathbb{Z}^{+}$ the user needed to answer $f_{v}$ correctly. We use the $\text{KAR}^{\text{3}}\text{L}$ Shu et al. (2024) model and flashcard learning interface to select the set of new flashcards to show users.
@@ 3.2.1 Expressed Preferences
For Likert ratings, if a user sees a mnemonic $m_{X}$ after answering card $f_{v}$ incorrectly, they rate $m_{X}$ on a 5-Likert scale (Figure 3). We call this preference label $y_{rate}$ and set $y_{rate}=\texttt{A}$ if the average Likert rating of $m_{A}$ is higher than $m_{B}$ across users (same for B and tie). For each $f_{v}$, users see only one of $m_{A}$ or $m_{B}$, so their rating cannot be biased by having already seen the other in the pair.
Figure 4: Screenshot of UI for pairwise comparisons.